Newton's law of universal gravitation gives the force between two objects of mass 1 and mass 2 separated by distance how as F=G(M1 M2)/R2 use dimensional analysis to find the unit of gravitational
1
Expert's answer
2021-07-15T13:16:04-0400

Gravitational force acting between two objects of masses m1 and m2 separated by distance r, 


"F = \frac{Gm_1m_2}{r^2}"


⟹ "G = \frac{Fr^2}{m_1m_2}"


Thus dimensional formula of G is "\frac{[MLT^{-2}][L^2]}{[M^2]}"


⟹ "G = M^{-1}L^3T^{-2}"